OpenStudy (andijo76):

HELP I am having a hard time coming up with an answer A store mixes Kentucky bluegrass worth $12 per pound and rye grass worth $16 the mixture is to sell for $13 per pound find how much of each should be used to make a 192 pound mixture how many #'s of Kentucky bluegrass should be in the mixture the answer i ended up with is 48 and that is not right

OpenStudy (mertsj):

Let x = the pounds of bluegrass Then 192-x= pounds of rye x pounds of bluegrass are worth 12x 192-x pounds of rye grass are worth 16(192-x) 192 pounds of the mixture are worth 13(192) Write the equation that says: value of bluegrass+ value of rye grass = value of mixture

OpenStudy (andijo76):

thanks i guess i will figure out the total pounds of kentuck bluegrass should be in the mixure only

OpenStudy (mertsj):

Did you write the equation?

OpenStudy (andijo76):

im trying to figure it out thanks lets say i have only had algebra for a few months in my whole 37 years lol so its a learning experience

OpenStudy (mertsj):

Did you read what I wrote? What is the value of the blue grass?

OpenStudy (andijo76):

i came up with 47 bags ugh i dont what i am doing wrong

OpenStudy (mertsj):

12x+16(192-x)=13(192) 12x+3072-16x=2496 -4x=-576 x=144 pounds of bluegrass 192-144=48 pounds of ryegrass

OpenStudy (andijo76):

oh ok i was working on this and i came up with the same answer thanks